# Spacetime Quasicrystals

### v1

@@ Abstract
Self-similar quasicrystals (like the famous Penrose and Ammann-Beenker tilings) are exceptional geometric structures in which long-range order, quasiperiodicity, non-crystallographic orientational symmetry, and discrete scale invariance are tightly interwoven in a beautiful way. In this paper, we show how such structures may be generalized from Euclidean space to Minkowski spacetime. We construct the first examples of such Lorentzian quasicrystals (the spacetime analogues of the Penrose or Ammann-Beenker tilings), and point out key novel features of these structures (compared to their Euclidean cousins). We end with some (speculative) ideas about how such spacetime quasicrystals might relate to reality. This includes an intriguing scenario in which our infinite (3+1)(3+1)D universe is embedded (like one of our spacetime quasicrystal examples) in a particularly symmetric (9+1)(9+1)D torus $T^{9,1}$ (which was previously found to yield the most symmetric toroidal compactification of the superstring). We suggest how this picture might help explain the mysterious seesaw relationship $M_{\rm Pl}M_{\rm vac}\approx M_{\rm EW}^{2}$ between the Planck, vacuum energy, and electroweak scales ($M_{\rm Pl}$, $M_{\rm vac}$, $M_{\rm EW}$).
@@ I Introduction
Figure 1: Left, a patch of the Penrose tiling [58] given by rhombs in purple, underlaid by its “inflation” in pink. Right top, the thin (red) and thick (blue) Penrose rhombs, along with their edge-matching rules and inflation rules. Right bottom, Conway’s version of the tiles: “darts” (red) and “kites” (blue) are given with their inflation rules. The two versions are equivalent since tilings built from kites and darts are mutually locally derivable from tilings by rhombs. For an introduction to the Penrose tiling, see Ref. [31].
@@
Motivated both by these physical connections, and also by pure mathematical interest, in this paper we show how self-similar quasicrystals may be naturally generalized from Euclidean space to Minkowski spacetime; and we construct the first examples of such Lorentzian quasicrystals (essentially the spacetime analogues of Euclidean quasicrystals like the Penrose tiling or the Ammann-Beenker tiling). [^1] We draw the reader’s attention to previous references [29, 65, 30, 38] that have studied dynamical/physical models that they dub “time quasicrsystals” or “spacetime quasicrystals.” These are something different from the spacetime patterns constructed in present paper, in which quasiperiodicity, non-crystallographic orientational symmetry, and discrete self-similarity are intrinsically knit together as in the Penrose or Ammann-Beenker tilings. We point out key novel features of these objects (compared to their Euclidean cousins), and end with some (speculative) ideas about how such spacetime quasicrystals might relate to reality.
@@
In Sec. II we introduce some background about lattices and their symmetries, Coxeter’s general theory of reflection groups, and algebraic number theory.
@@
In Sec. III, as a warm-up for spacetime, we first develop the theory of quasicrystals in ordinary Euclidean space. After reviewing the standard “cut-and-project” (C&P) scheme (for constructing a quasicrystal by taking an irrationally-sloped “cut” through a higher-dimensional lattice, and projecting nearby lattice points onto the cut), we describe a variant of particular interest, which we call the “symmetric cut-and-project” (sC&P) scheme, in which quasiperiodicity, non-crystallographic symmetry, and discrete scale-invariance are intrinsically interwoven in the desired fashion (as in the Penrose or Ammann-Beenker tilings). Whereas the most standard C&P scheme uses a “window” to choose lattice points near the cut, we also discuss a more general C&P scheme based on a “weighting function” (which weights different points differently, according to their distance from the cut), and this allows us to define “self-dual” and “globally-scale-invariant” quasicrystals. (This will be important when we turn to Lorentzian quasicrystals.) To illustrate all of this, we use sC&P to construct the standard Ammann-Beenker tiling, as well as its self-dual variant (see Fig. 7). We end by discussing the types of (quasi)symmetries/dualities that sC&P quasicrystals have, and delineating two distinct types of scale invariance (“local” and “global”).
@@
This section is mainly a review, although the reader may find certain aspects of our formulation novel or helpful, as it has been designed (i) to clarify the logical connection between quasiperiodicity, non-crystallographic symmetry and discrete self-similarity or scale-invariance; and (ii) relatedly, to make the generalization from Euclidean to Lorentzian spacetime simple and natural. (And, to our knowledge, our notions of self-dual and globally-scale-invariant quasicrystals are also new.)
@@
In Sec. IV we are finally ready to extend our discussion to spacetime. First, in Sec. IV.1, we introduce Lorentzian lattices (i.e. lattices in Minkowski spacetime) and their reflection symmetries, focusing on the key novel features compared to lattices in Euclidean space. Then in Sec. IV.2 we explain how the sC&P algorithm may be naturally extended to spacetime. Crucially, the quasicrystal’s non-crystallographic orientational symmetry is now a dramatically larger (in fact infinite!) group, and consequently, the sC&P window (or weighting function) must also be much more symmetric. In Sec. IV.3 we discuss further consequences of this fact: spacetime quasicrystals are a symmetric point set without an associated symmetric tiling; and they do not have local scale invariance, but instead can have global scale invariance or self-duality. Then, in Sec. IV.4, we construct the first two examples of such spacetime quasicrystals: (i) the $1+1$ dimensional (self-dual or globally-scale-invariant) spacetime quasicrystal obtained from the $(3+1)$ dimensional odd self-dual Lorentzian lattice $\textrm{I}_{3,1}$, and (ii) the four particularly interesting $3+1$ dimensional (self-dual or globally-scale-invariant) spacetime quasicrystals obtained from the remarkable $(9+1)$ dimensional even self-dual Lorentzian lattice $\textrm{II}_{9,1}$.
@@
In Sec. V we present some (speculative) ideas about how such spacetime quasicrystals might relate to reality.
@@
Sec. V.1 is devoted to one such speculation: In Ref. [52], Moore pointed out that the most symmetric toroidal compactification of the superstring was obtained by curling all 10 dimensions of spacetime (including time) into a particularly symmetric torus $T^{9,1}$ obtained by modding out (9+1)(9+1)D Minkowski space by the remarkable (9+1)(9+1)D even self-dual lattice $\mathbb{II}_{9,1}$: $T^{9,1}=\mathbb{R}^{9,1}/\mathbb{II}_{9,1}$. At the time, it seemed that this 10D torus couldn’t have anything to do with the real world and was of purely academic interest (as it seemed to have closed time like curves, and lack our $3+1$ macroscopic dimensions). However, we explain how the quasicrystal construction in Sec. IV.4.2 suggests a potential way around this conclusion: the most symmetric embeddings of (3+1)(3+1)D Minkowski spacetime $\mathbb{R}^{3,1}$ in the 10D torus $T^{9,1}$ are irrationally-sloped with respect to the torus, so that $\mathbb{R}^{3,1}$ wraps around and around the torus, densely filling it, without ever intersecting itself. An observer confined to this $\mathbb{R}^{3,1}$ would regard themselves as living in infinite 4D Minkowski spacetime, despite the fact that this spacetime is embedded in the completely compactified 10D spacetime torus $T^{9,1}$ (the universe in a nutshell!). Moreover, we explain some (very tentative) arguments suggesting that this scenario might even be a helpful starting point for explaining the mysteriously large hierarchies between the Planck scale $M_{\rm Pl}$, the electroweak scale $M_{\rm EW}$ and the vacuum energy scale $M_{\rm vac}$ ($M_{\rm Pl}\gg M_{\rm EW}\gg M_{\rm vac}$), as well as the seesaw relationship they are observed to obey ($M_{\rm Pl}M_{\rm vac}\approx M_{\rm EW}^{2}$).
@@ II.1 Lattices
We will regard $\mathbb{R}^{d}$ as a $d$-dimensional vector space $V$ equipped with a real-valued inner product $\langle\cdot,\cdot\rangle$. [^2] In Sec. IV, in order to pass from space to spacetime, we will generalize $\langle\cdot,\cdot\rangle$ from a (positive definite) inner product to a (indefinite) nondegenerate bilinear form. Since the inner product is positive definite, we can choose an orthonormal basis (or “Cartesian basis”) $\{{\bf e}_{\alpha}\}$ for $V$ with
@@ II.2 Coxeter Theory
Consider the vector space $V=\mathbb{R}^{d}$ equipped with the (positive definite) inner product $\langle\cdot,\cdot\rangle$. In the basis $\{{\bf v}_{\mu}\}$ ($\mu=1,\ldots,d$), the metric is a symmetric, positive-definite matrix $g_{\mu\nu}=\langle{\bf v}_{\mu}|{\bf v}_{\nu}\rangle$.
@@
A Coxeter group $G$ is an abstraction and generalization of a reflection group: it is a group generated by $n$ generators $S=\{R_{1},\ldots,R_{n}\}$, subject to the relations $(R_{i}R_{j})^{m_{ij}}=1$, where $m_{ij}$ is a symmetric integer $n\times n$ matrix $(m_{ij}=m_{ji},m_{ij}\in\mathbb{Z})$, with $m_{ii}=1$ along the diagonal, and $m_{ij}>1$ off the diagonal (i.e. when $i\neq j$). The pair $(G,S)$ is called a Coxeter system. The matrix $m_{ij}$ is called the Coxeter matrix, and the corresponding reflections $R_{i}$, mirrors $M_{i}$, and roots ${\bf r}_{i}$ ($i=1,\ldots,n)$ are called the fundamental reflections, fundamental mirrors, and fundamental roots, respectively. (The number $n$ of fundamental roots is called the rank of $G$.)
@@
If we draw all the mirrors in a given reflection group $G$, they divide up the space $Y^{k}$ into a set of congruent convex fundamental domains, which together form a tessellation $T$ of $Y^{k}$. If we pick one such fundamental domain (or Coxeter polytope) $\Gamma$, the fundamental roots are the outward pointing normal vectors to its bounding mirror “faces” (or the hyperplanes defining them). So the corresponding Coxeter diagram specifies the shape of the fundamental domain $\Gamma$: the number of its faces, and the angle at which each pair of faces meets (see Fig. 3). And the symmetries of the Coxeter diagram reflect symmetries of the Coxeter polytope itself: e.g. the $\mathbb{Z}_{2}$ symmetry of the $A_{n}$, $D_{n}$ and $E_{6}$ graphs, or the larger $S_{3}$ symmetry of the $D_{4}$ graph which gives rise to the famous triality symmetry of the $so(8)$ Lie algebra. Given a choice of fundamental domain $\Gamma$, successive reflections through its faces can map it to any other copy of $\Gamma$ in the tessellation $T$ of $Y^{k}$, so that the “tiles” in $T$ are in one-to-one correspondence with the elements in the group $G$.
The spherical Coxeter groups have been completely classified (see the left panel of Fig. 2); and, moreover, Coxeter showed that these are precisely the finite Coxeter groups. The affine Coxeter groups have also been completely classified (see the right panel of Fig. 2). The hyperbolic Coxeter groups have not been classified; although the hyperbolic reflection groups whose fundamental domains are simplices have been classified (see Secs. 6.8 & 6.9 and Figs. 2 & 3 in [40]). Vinberg found additional hyperbolic reflection groups (with non-simplicial fundamental domains) by studying the reflection symmetries of Lorentzian lattices (see Sec. IV and App. A).
@@
The crystallographic finite/spherical groups are closely related to affine groups: each affine diagram may be obtained by adding a single “extending node” to one of the crystallographic finite groups (see Fig. 2). The affine diagrams describe the full reflection symmetries of some periodic pattern in Euclidean space $\mathbb{R}^{n}$, while the crystallographic finite/spherical diagrams arise as the point subgroups of those periodic patterns that fix a point where n1n-1 of the mirrors cross. [^3] The crystallographic finite Coxeter groups (and their associated Coxeter-Dynkin diagrams) also play a fundamental role in the classification of the finite-dimensional semi-simple Lie algebras (and their associated Lie groups) [41]. Thus, the list of finite crystallographic Coxeter-Dinkin diagrams in Fig. 2 (namely $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $E_{6}$, $E_{7}$, $E_{8}$, $F_{4}$ and $G_{2}$) is also famous as the list of simple finite-dimensional Lie algebras. More general crystallographic Coxeter-Dynkin diagrams (including affine, hyperbolic and indefinite diagrams) play an analogous role in the classification of Kac-Moody algebras [43].
@@
By contrast, the non-crystallographic finite/spherical groups are those that do not arise as the point groups within a larger periodic pattern. From Fig. 2, the only finite irreducible noncrystallographic Coxeter groups occur in two-dimensions ($I_{2}^{n}$, the symmetries of a regular $n$-gon, with $n\neq 2,3,4,6$), three-dimensions ($H_{3}$, the symmetries of the icosahedron) or four-dimensions ($H_{4}$, the symmetries of the 600-cell, the four-dimensional analogue of the icosahedron). These non-crystallographic groups play an important role in the theory of quasiperiodic tilings and quasicrystals, where they govern the orientational symmetries of such structures: e.g. the 2D Penrose tiling (with decagonal orientational symmetry $I_{2}^{10}$) [58, 32], the 2D Ammann-Beenker tiling (with octagonal symmetry $I_{2}^{8}$) [2, 10], various 3D quasicrystals and tilings (with icosahedral symmetry $H_{3}$) [5], and the 4D Elser-Sloane quasicrystal (with symmetry $H_{4}$) [27].
@@
One way to obtain a non-crystallographic Coxeter subgroup $H$ of a Coxeter group $G$ is by folding its Dynkin diagram. This is process where we partition the fundamental generators of $G$ into subsets, where all the generators in a given subset commute with one another; and then, for each subset, we take the product of all the generators in that subset, and regard that product as a new fundamental generator. For example (see the upper left of Fig. 4), if we partition the 8 generators $\{R_{1},\ldots,R_{8}\}$ of $E_{8}$ into four subsets; $\{R_{4},R_{8}\}$, $\{R_{3},R_{5}\}$, $\{R_{2},R_{6}\}$, $\{R_{1},R_{7}\}$, and use these subsets to define the new generators $P_{1}=R_{4}R_{8}$, $P_{2}=R_{3}R_{5}$, $P_{3}=R_{2}R_{6}$ and $P_{4}=R_{1}R_{7}$, we find that the $P_{i}$’s obey the Coxeter relations of non-crystallographic Coxeter subgroup $H_{4}\in E_{8}$. Fig. 4 illustrates some important examples in which finite Coxeter groups are folded to obtain their non-crystallographic subgroups. (These are the most important foldings from the standpoint of constructing ordinary Euclidean quasicrystals.)

### v2

@@ Abstract
Self-similar quasicrystals (like the famous Penrose and Ammann-Beenker tilings) are exceptional geometric structures in which long-range order, quasi-periodicity, non-crystallographic orientational symmetry, and discrete scale invariance are tightly interwoven in a beautiful way. In this paper, we show how such structures may be generalized from Euclidean space to Minkowski spacetime. We construct the first examples of such Lorentzian quasicrystals (the spacetime analogues of the Penrose or Ammann-Beenker tilings), and point out key novel features of these structures (compared to their Euclidean cousins). We end with some (speculative) ideas about how such spacetime quasicrystals might relate to reality. This includes an intriguing scenario in which our infinite (3+1)(3+1)D universe is embedded (like one of our spacetime quasicrystal examples) in a particularly symmetric (9+1)(9+1)D torus $T^{9,1}$ (which was previously found to yield the most symmetric toroidal compactification of the superstring). We suggest how this picture might help explain the mysterious seesaw relationship $M_{\rm Pl}M_{\rm vac}\approx M_{\rm EW}^{2}$ between the Planck, vacuum energy, and electroweak scales ($M_{\rm Pl}$, $M_{\rm vac}$, $M_{\rm EW}$).
@@ I Introduction
Figure 1: Left: a patch of the Penrose tiling [59] given by rhombs in purple, underlaid by its “inflation” in pink. Right top: the thin (red) and thick (blue) Penrose rhombs, along with their edge-matching rules and inflation rules. Right bottom: Conway’s version of the tiles: “darts” (red) and “kites” (blue) are given with their inflation rules. The two versions are equivalent since tilings built from kites and darts are mutually locally derivable from tilings by rhombs. For an introduction to the Penrose tiling, see Ref. [32].
@@
Motivated both by these physical connections, and also by pure mathematical interest, in this paper we show how self-similar quasicrystals may be naturally generalized from Euclidean space to Minkowski spacetime; and we construct the first examples of such Lorentzian quasicrystals (essentially the spacetime analogues of Euclidean quasicrystals like the Penrose tiling or the Ammann-Beenker tiling). [^1] We draw the reader’s attention to previous references [30, 67, 31, 39] that have studied dynamical/physical models that they dub “time quasicrystals” or “spacetime quasicrystals.” These are something different from the spacetime patterns constructed in present paper, in which quasi-periodicity, non-crystallographic orientational symmetry, and discrete self-similarity are intrinsically knit together as in the Penrose or Ammann-Beenker tilings. We point out key novel features of these objects (compared to their Euclidean cousins), and end with some (speculative) ideas about how such spacetime quasicrystals might relate to reality.
@@
In Sec. II, we introduce some background about lattices and their symmetries, Coxeter’s general theory of reflection groups, and algebraic number theory.
@@
In Sec. III, as a warm-up for spacetime, we first develop the theory of quasicrystals in ordinary Euclidean space. After reviewing the standard “cut-and-project” (C&P) scheme (for constructing a quasicrystal by taking an irrationally-sloped “cut” through a higher-dimensional lattice, and projecting nearby lattice points onto the cut), we describe a variant of particular interest, which we call the “symmetric cut-and-project” (sC&P) scheme, in which quasi-periodicity, non-crystallographic symmetry, and discrete scale-invariance are intrinsically interwoven in the desired fashion (as in the Penrose or Ammann-Beenker tilings). Whereas the most standard C&P scheme uses a “window” to choose lattice points near the cut, we also discuss a more general C&P scheme based on a “weighting function” (which weights different points differently, according to their distance from the cut), and this allows us to define “self-dual” and “globally-scale-invariant” quasicrystals (this will be important when we turn to Lorentzian quasicrystals.) To illustrate all of this, we use sC&P to construct the standard Ammann-Beenker tiling, as well as its self-dual variant (see Fig. 7). We end by discussing the types of (quasi)symmetries/dualities that sC&P quasicrystals have, and delineating two distinct types of scale invariance (“local” and “global”).
@@
This section is mainly a review, although the reader may find certain aspects of our formulation novel or helpful, as it has been designed (i) to clarify the logical connection between quasi-periodicity, non-crystallographic symmetry and discrete self-similarity or scale-invariance; and (ii) relatedly, to make the generalization from Euclidean to Lorentzian spacetime simple and natural (and, to our knowledge, our notions of self-dual and globally-scale-invariant quasicrystals are also new.)
@@
In Sec. IV, we are finally ready to extend our discussion to spacetime. First, in Sec. IV.1, we introduce Lorentzian lattices (i.e. lattices in Minkowski spacetime) and their reflection symmetries, focusing on the key novel features compared to lattices in Euclidean space. Then in Sec. IV.2, we explain how the sC&P algorithm may be naturally extended to spacetime. Crucially, the quasicrystal’s non-crystallographic orientational symmetry is now a dramatically larger (in fact infinite!) group, and consequently, the sC&P window (or weighting function) must also be much more symmetric. In Sec. IV.3, we discuss further consequences of this fact: spacetime quasicrystals are a symmetric point set without an associated symmetric tiling; and they do not have local scale invariance, but instead can have global scale invariance or self-duality. Then, in Sec. IV.4, we construct the first two examples of such spacetime quasicrystals: (i) the $1+1$ dimensional (self-dual or globally-scale-invariant) spacetime quasicrystal obtained from the $(3+1)$ dimensional odd self-dual Lorentzian lattice $\textrm{I}_{3,1}$, and (ii) the four particularly interesting $3+1$ dimensional (self-dual or globally-scale-invariant) spacetime quasicrystals obtained from the remarkable $(9+1)$ dimensional even self-dual Lorentzian lattice $\textrm{II}_{9,1}$.
@@
In Sec. V, we present some (speculative) ideas about how such spacetime quasicrystals might relate to reality.
@@
Sec. V.1 is devoted to one such speculation: In Ref. [53], Moore pointed out that the most symmetric toroidal compactification of the superstring was obtained by curling all 10 dimensions of spacetime (including time) into a particularly symmetric torus $T^{9,1}$ obtained by modding out (9+1)(9+1)D Minkowski space by the remarkable (9+1)(9+1)D even self-dual lattice $\mathrm{II}_{9,1}$: $T^{9,1}=\mathbb{R}^{9,1}/\mathrm{II}_{9,1}$. At the time, it seemed that this 10D torus couldn’t have anything to do with the real world and was of purely academic interest (as it seemed to have closed time like curves, and lack our $3+1$ macroscopic dimensions). However, we explain how the quasicrystal construction in Sec. IV.4.2 suggests a potential way around this conclusion: the most symmetric embeddings of (3+1)(3+1)D Minkowski spacetime $\mathbb{R}^{3,1}$ in the 10D torus $T^{9,1}$ are irrationally-sloped with respect to the torus, so that $\mathbb{R}^{3,1}$ wraps around and around the torus, densely filling it, without ever intersecting itself. An observer confined to this $\mathbb{R}^{3,1}$ would regard themselves as living in infinite 4D Minkowski spacetime, despite the fact that this spacetime is embedded in the completely compactified 10D spacetime torus $T^{9,1}$ (the universe in a nutshell!). Moreover, we explain some (very tentative) arguments suggesting that this scenario might even be a helpful starting point for explaining the mysteriously large hierarchies between the Planck scale $M_{\rm Pl}$, the electroweak scale $M_{\rm EW}$ and the vacuum energy scale $M_{\rm vac}$ ($M_{\rm Pl}\gg M_{\rm EW}\gg M_{\rm vac}$), as well as the seesaw relationship they are observed to obey ($M_{\rm Pl}M_{\rm vac}\approx M_{\rm EW}^{2}$).
@@ II.1 Lattices
We will regard $\mathbb{R}^{d}$ as a $d$-dimensional vector space $V$ equipped with a real-valued inner product $\langle\cdot,\cdot\rangle$. [^2] In Sec. IV, in order to pass from space to spacetime, we will generalize $\langle\cdot,\cdot\rangle$ from a (positive definite) inner product to a (indefinite) non-degenerate bilinear form. Since the inner product is positive definite, we can choose an orthonormal basis (or “Cartesian basis”) $\{{\bf e}_{\alpha}\}$ for $V$ with
@@ II.2 Coxeter Theory
Consider the vector space $V=\mathbb{R}^{d}$ equipped with the (positive-definite) inner product $\langle\cdot,\cdot\rangle$. In the basis $\{{\bf v}_{\mu}\}$ ($\mu=1,\ldots,d$), the metric is a symmetric, positive-definite matrix $g_{\mu\nu}=\langle{\bf v}_{\mu}|{\bf v}_{\nu}\rangle$.
@@
A Coxeter group $G$ is an abstraction and generalization of a reflection group: it is a group generated by $n$ generators $S=\{R_{1},\ldots,R_{n}\}$, subject to the relations $(R_{i}R_{j})^{m_{ij}}=1$, where $m_{ij}$ is a symmetric integer $n\times n$ matrix $(m_{ij}=m_{ji},m_{ij}\in\mathbb{Z})$, with $m_{ii}=1$ along the diagonal, and $m_{ij}>1$ off the diagonal (i.e. when $i\neq j$). The pair $(G,S)$ is called a Coxeter system. The matrix $m_{ij}$ is called the Coxeter matrix, and the corresponding reflections $R_{i}$, mirrors $M_{i}$, and roots ${\bf r}_{i}$ ($i=1,\ldots,n)$ are called the fundamental reflections, fundamental mirrors, and fundamental roots, respectively. The number $n$ of fundamental roots is called the rank of $G$.
@@
If we draw all the mirrors in a given reflection group $G$, they divide up the space $Y^{k}$ into a set of congruent convex fundamental domains, which together form a tessellation $T$ of $Y^{k}$. If we pick one such fundamental domain (or Coxeter polytope) $\Gamma$, the fundamental roots are the outward pointing normal vectors to its bounding mirror “faces” (or the hyperplanes defining them). So the corresponding Coxeter diagram specifies the shape of the fundamental domain $\Gamma$: the number of its faces, and the angle at which each pair of faces meets (see Fig. 3). And the symmetries of the Coxeter diagram reflect symmetries of the Coxeter polytope itself: e.g. the $\mathbb{Z}_{2}$ symmetry of the $A_{n}$, $D_{n}$ and $E_{6}$ graphs, or the larger $S_{3}$ symmetry of the $D_{4}$ graph which gives rise to the famous triality symmetry of the $\mathfrak{so}(8)$ Lie algebra. Given a choice of fundamental domain $\Gamma$, successive reflections through its faces can map it to any other copy of $\Gamma$ in the tessellation $T$ of $Y^{k}$, so that the “tiles” in $T$ are in one-to-one correspondence with the elements in the group $G$.
The spherical Coxeter groups have been completely classified (see the left panel of Fig. 2); and, moreover, Coxeter showed that these are precisely the finite Coxeter groups. The affine Coxeter groups have also been completely classified (see the right panel of Fig. 2). The hyperbolic Coxeter groups have not been classified, although the hyperbolic reflection groups whose fundamental domains are simplices have been classified (see Secs. 6.8 & 6.9 and Figs. 2 & 3 in [41]). Vinberg found additional hyperbolic reflection groups (with non-simplicial fundamental domains) by studying the reflection symmetries of Lorentzian lattices (see Sec. IV and App. A).
@@
The crystallographic finite/spherical groups are closely related to affine groups: each affine diagram may be obtained by adding a single “extending node” to one of the crystallographic finite groups (see Fig. 2). The affine diagrams describe the full reflection symmetries of some periodic pattern in Euclidean space $\mathbb{R}^{n}$, while the crystallographic finite/spherical diagrams arise as the point subgroups of those periodic patterns that fix a point where n1n-1 of the mirrors cross. [^3] The crystallographic finite Coxeter groups (and their associated Coxeter-Dynkin diagrams) also play a fundamental role in the classification of the finite-dimensional semi-simple Lie algebras (and their associated Lie groups) [42]. Thus, the list of finite crystallographic Coxeter-Dynkin diagrams in Fig. 2 (namely $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $E_{6}$, $E_{7}$, $E_{8}$, $F_{4}$ and $G_{2}$) is also famous as the list of simple finite-dimensional Lie algebras. More general crystallographic Coxeter-Dynkin diagrams (including affine, hyperbolic and indefinite diagrams) play an analogous role in the classification of Kac-Moody algebras [44].
@@
By contrast, the non-crystallographic finite/spherical groups are those that do not arise as the point groups within a larger periodic pattern. From Fig. 2, the only finite irreducible non-crystallographic Coxeter groups occur in two dimensions ($I_{2}^{n}$, the symmetries of a regular $n$-gon, with $n\neq 2,3,4,6$), three dimensions ($H_{3}$, the symmetries of the icosahedron) or four dimensions ($H_{4}$, the symmetries of the 600-cell, the four-dimensional analogue of the icosahedron). These non-crystallographic groups play an important role in the theory of quasi-periodic tilings and quasicrystals, where they govern the orientational symmetries of such structures: e.g. the 2D Penrose tiling (with decagonal orientational symmetry $I_{2}^{10}$) [59, 33], the 2D Ammann-Beenker tiling (with octagonal symmetry $I_{2}^{8}$) [2, 10], various 3D quasicrystals and tilings (with icosahedral symmetry $H_{3}$) [5], and the 4D Elser-Sloane quasicrystal (with symmetry $H_{4}$) [28].
@@
One way to obtain a non-crystallographic Coxeter subgroup $H$ of a Coxeter group $G$ is by folding its Dynkin diagram. This is a process where we partition the fundamental generators of $G$ into subsets, where all the generators in a given subset commute with one another; and then, for each subset, we take the product of all the generators in that subset, and regard that product as a new fundamental generator (see also [26]). For example (see the upper left of Fig. 4), if we partition the 8 generators $\{R_{1},\ldots,R_{8}\}$ of $E_{8}$ into four subsets; $\{R_{4},R_{8}\}$, $\{R_{3},R_{5}\}$, $\{R_{2},R_{6}\}$, $\{R_{1},R_{7}\}$, and use these subsets to define the new generators $\rho_{1}=R_{4}R_{8}$, $\rho_{2}=R_{3}R_{5}$, $\rho_{3}=R_{2}R_{6}$ and $\rho_{4}=R_{1}R_{7}$, we find that the $\rho_{i}$’s obey the Coxeter relations of non-crystallographic Coxeter subgroup $H_{4}\in E_{8}$. Fig. 4 illustrates some important examples in which finite Coxeter groups are folded to obtain their non-crystallographic subgroups. These are the most important foldings from the standpoint of constructing ordinary Euclidean quasicrystals. All foldings that lead to the 2D non-crystallographic $I_{2}^{n}$ arise from grouping all blue and red generators together to form $\rho_{B}$ and $\rho_{R}$, and correspond geometrically to a projection onto the 2D Coxeter planes (see Subsections II.4 and IV.1.3).